OpenStudy (jazzyfa30):

2^3 - 4[3^2 -2(9^2 - 79)]

OpenStudy (anonymous):

=>8-4[9-2(81-79)] =>8-4[9-2*2] =>8-4[9-4] =>8-4*5 =8-20 =-12 Ans

OpenStudy (jazzyfa30):

thank you